Calculate the Velocity when time is 1.5seconds, the Velocity equation is V=30m/s+(25m/s³)t²
1
Expert's answer
2020-05-05T18:39:10-0400

"V=30+25t^2\\V=30+25\times1.5\times1.5=86.25m/s"